I don't know what I'm doing wrong. I triple checked with the 2nd question: Calculators online, my scientific calculator, everything comes up with the same number but per system, my answer is wrong. Please help(Related to Checkpoint 5.3) (Compound interest with non-annual periods) Calculate the amount of money that will be in each of
Answer #1

Calculating the Future Value of Deposit at the end of different periods:-

a).

Future Value = Invested Amount(1+\frac{r}{m})^{n*m}

Where, Invested Amount = $1000

r = Periodic Interest rate = 12%

n= no of periods = 10

m = no of times compounding in a year = 4

Future Value =1000*(1+\frac{0.12}{4})^{10*4}

Future Value =1000*3.262037792

Future Value = $3262.04

b).

Future Value = Invested Amount(1+\frac{r}{m})^{n*m}

Where, Invested Amount = $96,000

r = Periodic Interest rate = 10%

n= no of periods = 3

m = no of times compounding in a year = 2

Future Value =96000*(1+\frac{0.10}{2})^{2*3}

FutureValue 96000 *(1 + 0.05)

Future Value =96000*1.34009564063

Future Value = $128,649.18

c).

Future Value = Invested Amount(1+\frac{r}{m})^{n*m}

Where, Invested Amount = $9,000

r = Periodic Interest rate = 12%

n= no of periods = 6

m = no of times compounding in a year = 3

Future Value =9000*(1+\frac{0.12}{3})^{3*6}

Future Value =9000*(1+0.04)^{18}

Future Value =9000*2.02581651538

Future Value = $18,232.35

d).

Future Value = Invested Amount(1+\frac{r}{m})^{n*m}

Where, Invested Amount = $118,000

r = Periodic Interest rate = 8%

n= no of periods = 4

m = no of times compounding in a year = 6

Future Value =118000*(1+\frac{0.08}{6})^{6*4}

Future Value =118000*(1+0.01333)^{24}

Future Value =118000*1.37421882444

Future Value = $162,157.82

e).

Future Value = Invested Amount(1+\frac{r}{m})^{n*m}

Where, Invested Amount = $30,000

r = Periodic Interest rate = 18%

n= no of periods = 4

m = no of times compounding in a year = 12

Future Value =30000*(1+\frac{0.18}{12})^{4*12}

Future Value =30000*(1+0.015)^{48}

Future Value =30,000*2.04347828931

Future Value = $61,304.35

f).

Future Value = Invested Amount(1+\frac{r}{m})^{n*m}

Where, Invested Amount = $17,000

r = Periodic Interest rate = 10%

n= no of periods = 3

m = no of times compounding in a year = 1

Future Value =17000*(1+\frac{0.10}{1})^{3*1}

Future Value =17000*(1+0.10)^{3}

Future Value =17,000*1.331

Future Value = $22,627

If you need any clarification, you can ask in comments.     

If you like my answer, then please up-vote as it will be motivating